# Non-Convex Self-Concordant Functions: Practical Algorithms and Complexity Analysis

### v1

@@ 3.2 Adaptive Regularization Method (ARM)
$d_{j}=\begin{cases}-(\nabla^{2}(f+\sigma_{j}F)(x_{j}))^{-1}\nabla f(x_{j}),&\textnormal{if }\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}},\\ v_{j},&\textnormal{if }\nabla f(x_{j})^{\top}v_{j}\leq 0,\ \lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}},\\ -v_{j},&\textnormal{otherwise,}\end{cases}$ (17)
where $\lambda_{\mathrm{nc}}=\sigma_{j}\sqrt{\epsilon_{H}}\nabla^{2}F(x_{j})[v_{j},v_{j}]$. This rule ensures that, if the significant negative curvature is detected, a descent direction aligned with it will be used to escape saddle regions. Under the same notation of Proposition 2.6, we view $f$ as $\sigma_{j}F$-based $\kappa$-self-concordant and define the local model $m_{j}(t)$ as follows,
$\small m_{j}(t)=\begin{cases}f(x_{j})-t(\nu_{f,\sigma_{j}F}(x_{j}))^{2}+\kappa^{-2}\omega_{\star}(\kappa t\nu_{f,\sigma_{j}F}(x_{j})),&\text{if }\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}},\\ f(x_{j})-\kappa_{F}^{-2}\omega(\kappa_{F}t\sqrt{\sigma_{j}\Delta_{j}})+\kappa^{-2}\omega_{\star}(\kappa t\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}),&\text{otherwise.}\end{cases}$ (18)
If $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$, we have that $m_{j}(t)$ reduces to the one used for adaptive regularized Newton’s methods. Thus, we can use Eq. (16). If $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$ and $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$, Algorithm 1 performs a negative curvature step where $\delta_{j}=\lambda_{\min}(\nabla^{2}f(x_{j}))<0$. Minimizing $m_{j}(t)$ with respect to $t$ yields
$t_{j}=-\tfrac{\delta_{j}}{\sqrt{\sigma_{j}\Delta_{j}}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}(\kappa_{F}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}+\kappa\sqrt{\sigma_{j}\Delta_{j}})}.$ (19)
@@
Suppose that Assumption 3.1 holds with a $\kappa_{F}$-self-concordant $F$. Let $d_{j}$ and $m_{j}$ be defined in Eq. (17) and Eq. (18), and the termination condition be $\nu_{f,\sigma_{j}F}(x_{j})\leq\epsilon_{g}$ and $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\sigma_{j}\sqrt{\epsilon_{H}}\nabla^{2}F(x_{j})[v_{j},v_{j}]$. Then, we have $\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$ for all $j$, and Algorithm 1 terminates within at most $O(\epsilon_{g}^{-2}+\epsilon_{H}^{-1.5})$ iterations.
@@
The ARM improves the RNM in two dimensions. First, it introduces a parameter $\sigma_{j}$ that controls the magnitude of regularization, mitigating the issue of excessively large regularization, as discussed in Remark 3.4. Second, the ARM is effective when $f$ is $\ell F$-based $\kappa$-self-concordant, but we have no prior knowledge of $\ell$.

### v2

@@ 3.2 Adaptive Regularization Method (ARM)
$d_{j}=\begin{cases}-(\nabla^{2}(f+\sigma_{j}F)(x_{j}))^{+}\nabla f(x_{j}),&\textnormal{if }\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}},\\ v_{j},&\textnormal{if }\nabla f(x_{j})^{\top}v_{j}\leq 0,\ \lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}},\\ -v_{j},&\textnormal{otherwise,}\end{cases}$ (14)
where $\lambda_{\mathrm{nc}}=\sigma_{j}\sqrt{\epsilon_{H}}\nabla^{2}F(x_{j})[v_{j},v_{j}]$. This rule ensures that, if the significant negative curvature is detected, a descent direction aligned with it will be used to escape saddle regions. We define the local model $m_{j}(t)$ as follows,
$m_{j}(t)=\begin{cases}f(x_{j})-t(\nu_{f,\sigma_{j}F}(x_{j}))^{2}+\kappa^{-2}\omega_{\star}(\kappa t\nu_{f,\sigma_{j}F}(x_{j})),&\text{if }\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}},\\ f(x_{j})-\kappa_{F}^{-2}\omega(\kappa_{F}t\sqrt{\sigma_{j}\Delta_{j}})+\kappa^{-2}\omega_{\star}(\kappa t\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}),&\text{otherwise.}\end{cases}$ (15)
If $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$, we have that $m_{j}(t)$ reduces to the one used for adaptive regularized Newton’s methods (see Eq. (13)). Thus, we can use Eq. (13). If $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$ and $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$, Algorithm 2 performs a negative curvature step where $\delta_{j}=\lambda_{\min}(\nabla^{2}f(x_{j}))<0$. Minimizing $m_{j}(t)$ with respect to $t$ yields
$t_{j}=\begin{cases}-\tfrac{\delta_{j}}{\sqrt{\sigma_{j}\Delta_{j}}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}(\kappa_{F}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}+\kappa\sqrt{\sigma_{j}\Delta_{j}})},&\mathrm{if}\ \lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}},\\ \tfrac{1}{1+\kappa\nu_{f,\sigma_{j}F}(x_{j})},&\mathrm{if}\ \lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}.\end{cases}$ (16)
@@
Suppose that Assumption 3.1 holds with a $\kappa_{F}$-self-concordant and strictly convex $F$ and $d_{j}$ and $m_{j}$ are defined by Eqs. (14) and (15) for all $j$. Then, at all iterations of ARM with Opt =2=2, we have $\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$ for all $j$, and the option 2 variant of Algorithm 2 terminates within at most $O(\epsilon_{g}^{-2}+\epsilon_{H}^{-1.5})$ iterations.
Proof.
Proof We prove that $\sigma_{j}$ is uniformly upper bounded by contradiction. Assume there exists an index $j$ such that $\sigma_{j+1}>\max(\sigma_{0},\gamma_{3})\geq\sigma_{j}$. By the update rule for $\sigma$ in Algorithm 2, we always have $\sigma_{j+1}\leq\gamma_{3}\sigma_{j}$; hence $\sigma_{j}\geq\sigma_{j+1}/\gamma_{3}>1$. Assumption 3.1 implies $\delta_{j}+\Delta_{j}\geq 0$ and hence $\delta_{j}+\sigma_{j}\Delta_{j}\geq 0$.
We consider two cases. Case 1: $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$. Then $d_{j}$ is the regularized Newton direction and $m_{j}$ reduces to the Newton-type model. By Proposition 2.6, we have $x_{j}+t_{j}d_{j}\in\textbf{dom}(f)$ and
$f(x_{j})-f(x_{j}+t_{j}d_{j})\ \geq\ \rho_{j}t_{j}-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}\bigr).$
Comparing with the definition of $m_{j}$ in this case and using $\sigma_{j}\geq 1$ together with the monotonicity of $\omega_{\star}$ yields $f(x_{j})-f(x_{j}+t_{j}d_{j})\geq f(x_{j})-m_{j}(t_{j})$, i.e., $r_{j}\geq 1>\eta_{2}$. Algorithm 2 would therefore enforce $\sigma_{j+1}\leq\sigma_{j}$, contradicting $\sigma_{j+1}>\sigma_{j}$. Case 2: $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$. Then $d_{j}=\pm v_{j}$ and $\delta_{j}=\lambda_{\min}(\nabla^{2}f(x_{j}))<0$. In this case,
$f(x_{j})-m_{j}(t_{j})=-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\sigma_{j}\Delta_{j}}\bigr)+\kappa_{F}^{-2}\omega\!\bigl(\kappa_{F}t_{j}\sqrt{\sigma_{j}\Delta_{j}}\bigr).$
By Proposition 2.6, we also have
$f(x_{j})-f(x_{j}+t_{j}d_{j})\geq-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\Delta_{j}}\bigr)+\kappa_{F}^{-2}\omega\!\bigl(\kappa_{F}t_{j}\sqrt{\Delta_{j}}\bigr).$
We define $g(\sigma):=-\kappa^{-2}\omega_{\star}\!\bigl(\kappa t_{j}\sqrt{\delta_{j}+\sigma\Delta_{j}}\bigr)+\kappa_{F}^{-2}\omega\!\bigl(\kappa_{F}t_{j}\sqrt{\sigma\Delta_{j}}\bigr)$. A direct differentiation shows
$g^{\prime}(\sigma)=\tfrac{t_{j}^{2}\Delta_{j}}{2}\Bigl(-\tfrac{1}{1-\kappa t_{j}\sqrt{\delta_{j}+\sigma\Delta_{j}}}+\tfrac{1}{1+\kappa_{F}t_{j}\sqrt{\sigma\Delta_{j}}}\Bigr)\leq 0,$
and hence, $g(\sigma_{j})\leq g(1)$. Therefore $f(x_{j})-m_{j}(t_{j})=g(\sigma_{j})\leq g(1)\leq f(x_{j})-f(x_{j}+t_{j}d_{j})$, which implies $r_{j}\geq 1>\eta_{2}$ as in Case 1. Therefore, We conclude that $\sigma_{j}\leq\max(\sigma_{0},\gamma_{3})$ for all $j\geq 0$.
The argument below follows the usual successful/unsuccessful iteration counting used in adaptive regularization [11, 12] and trust-region analyses [60]; we include it in full for completeness. Fix an integer $k\geq 1$ and suppose Algorithm 2 has not terminated before iteration $k$. Define the sets of accepted (successful) and rejected (unsuccessful) iterations up to $k$ by
$\mathcal{S}_{k}:=\{0\leq j\leq k-1:\ r_{j}>\eta_{1}\},\qquad{\rm and}\qquad\mathcal{U}_{k}:=\{0\leq j\leq k-1:\ r_{j}\leq\eta_{1}\}.$
For any $j\in\mathcal{S}_{k}$, the step is accepted, so $x_{j+1}=x_{j}+t_{j}d_{j}$ and
$f(x_{j})-f(x_{j+1})=f(x_{j})-f(x_{j}+t_{j}d_{j})=r_{j}\bigl(f(x_{j})-m_{j}(t_{j})\bigr)\geq\eta_{1}\bigl(f(x_{j})-m_{j}(t_{j})\bigr).$
For any $j\in\mathcal{U}_{k}$, the step is rejected, so $x_{j+1}=x_{j}$ and $f(x_{j})-f(x_{j+1})=0$. Let $f_{\inf}:=\inf_{x\in\mathbb{R}^{n}}f(x)$ (finite under Assumption 3.1). Summing over $j=0,\dots,k-1$ yields
$f(x_{0})-f_{\inf}\geq\textstyle\sum_{j=0}^{k-1}f(x_{j})-f(x_{j+1})=\textstyle\sum_{j\in\mathcal{S}_{k}}f(x_{j})-f(x_{j+1})\geq\eta_{1}\textstyle\sum_{j\in\mathcal{S}_{k}}f(x_{j})-m_{j}(t_{j}).$ (17)
We now lower bound $f(x_{j})-m_{j}(t_{j})$ for $j\in\mathcal{S}_{k}$. Since the algorithm has not terminated at $x_{j}$, either (i) $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$ and $\nu_{f,\sigma_{j}F}(x_{j})>\epsilon_{g}$, or (ii) $\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}$. In case (i), using Eq. (13), we obtain $f(x_{j})-m_{j}(t_{j})\ \geq\ \kappa^{-2}\omega(\kappa\nu_{f,\sigma_{j}F}(x_{j}))\ \geq\ \kappa^{-2}\omega(\kappa\epsilon_{g})$.
In case (ii), set $z:=-\delta_{j}/(\sigma_{j}\Delta_{j})$. Since $\delta_{j}=\lambda_{\min}(\nabla^{2}f(x_{j}))<-\lambda_{\mathrm{nc}}=-\sigma_{j}\sqrt{\epsilon_{H}}\Delta_{j}$, we have $z>\sqrt{\epsilon_{H}}$ and, for the step size $t_{j}$ in Eq. (16), the explicit form of $m_{j}$ yields $f(x_{j})-m_{j}(t_{j})\geq\tfrac{z^{3}}{6(\kappa+\kappa_{F})^{2}}\ \geq\ \tfrac{\epsilon_{H}^{1.5}}{6(\kappa+\kappa_{F})^{2}}$. Consequently, for all $j\in\mathcal{S}_{k}$, $f(x_{j})-m_{j}(t_{j})\ \geq\ \widetilde{\epsilon}:=\min\!\left(\kappa^{-2}\omega(\kappa\epsilon_{g}),\ \tfrac{\epsilon_{H}^{1.5}}{6(\kappa+\kappa_{F})^{2}}\right)$. Plugging this into Eq. (17) gives $f(x_{0})-f_{\inf}\ \geq\ \eta_{1}\,\lvert\mathcal{S}_{k}\rvert\,\widetilde{\epsilon}$, and hence $\lvert\mathcal{S}_{k}\rvert\ \leq\ \frac{f(x_{0})-f_{\inf}}{\eta_{1}\,\widetilde{\epsilon}}=O(\widetilde{\epsilon}^{-1})$.
It remains to bound the number of rejected iterations $\lvert\mathcal{U}_{k}\rvert$. For any $j\in\mathcal{S}_{k}$ we have $\sigma_{j+1}\geq\max(\sigma_{\min},\gamma_{1}\sigma_{j})\geq\gamma_{1}\sigma_{j}$, while for any $j\in\mathcal{U}_{k}$ we have $\sigma_{j+1}\geq\gamma_{2}\sigma_{j}$. Thus, letting $\sigma_{\max}=\max(\sigma_{0},\gamma_{3})$, we have $\sigma_{0}\,\gamma_{1}^{\lvert\mathcal{S}_{k}\rvert}\gamma_{2}^{\lvert\mathcal{U}_{k}\rvert}\ \leq\ \sigma_{k}\ \leq\ \sigma_{\max}$. Taking logarithms and rearranging yields $\lvert\mathcal{U}_{k}\rvert\leq\tfrac{\log(\sigma_{\max}/\sigma_{0})-\lvert\mathcal{S}_{k}\rvert\log(\gamma_{1})}{\log(\gamma_{2})}$, and therefore $k=\lvert\mathcal{S}_{k}\rvert+\lvert\mathcal{U}_{k}\rvert=O(\lvert\mathcal{S}_{k}\rvert)=O(\widetilde{\epsilon}^{-1})$. Finally, since
$\widetilde{\epsilon}^{-1}\leq\kappa^{2}(\omega(\kappa\epsilon_{g}))^{-1}+\tfrac{6(\kappa+\kappa_{F})^{2}}{\epsilon_{H}^{1.5}}=O(\epsilon_{g}^{-2}+\epsilon_{H}^{-1.5}),$
the claimed iteration bound follows. $\square$
@@
Suppose that Algorithm 2 generates a nonzero trial step, i.e., $t_{j}d_{j}\neq 0$. Then $\delta_{j}+\sigma_{j}\Delta_{j}>0$. Indeed, by the convention $\sqrt{x}=+\infty$ for $x<0$ and the rule in Line 6 of Algorithm 2 that sets $t_{j}=0$ whenever the minimizer of $m_{j}$ is attained at $\infty$ or $m_{j}$ is constant, a nonzero trial step can occur only when the model $m_{j}$ admits a finite minimizer at some $t_{j}>0$, which excludes the cases $\delta_{j}+\sigma_{j}\Delta_{j}\leq 0$. Moreover, if Option 1 is used, or if Option 2 is used with $\lambda_{\min}(\nabla^{2}f(x_{j}))\geq-\lambda_{\mathrm{nc}}$, then $\rho_{j}>0$. In the negative curvature branch of Option 2, i.e., Line 2 of Eq. (15), we only require $\rho_{j}\geq 0$, which is achieved by Eq. (14).
@@
ARM improves on RNM in two ways: First, ARM introduces a parameter $\sigma_{j}$ that controls the magnitude of regularization, mitigating the issue of excessively large regularization. Second, ARM is effective when $f$ is $\ell F$-based $\kappa$-self-concordant, but we have no prior knowledge of $\ell$.
Remark 3.10.